# Mixture of Experts based Multi-task Supervise Learning from Crowds

### v1

@@ Abstract
Existing truth inference methods in crowdsourcing aim to map redundant labels and items to the ground truth. They treat the ground truth as hidden variables and use statistical or deep learning-based worker behavior models to infer the ground truth. However, worker behavior models that rely on ground truth hidden variables overlook workers’ behavior at the item feature level, leading to imprecise characterizations and negatively impacting the quality of truth inference. This paper proposes a new paradigm of multi-task supervised learning from crowds, which eliminates the need for modeling of items’s ground truth in worker behavior models. Within this paradigm, we propose a worker behavior model at the item feature level called Mixture of Experts based Multi-task Supervised Learning from Crowds (MMLC). Two truth inference strategies are proposed within MMLC. The first strategy, named MMLC-owf, utilizes clustering methods in the worker spectral space to identify the projection vector of the oracle worker. Subsequently, the labels generated based on this vector are considered as the inferred truth. The second strategy, called MMLC-df, employs the MMLC model to fill the crowdsourced data, which can enhance the effectiveness of existing truth inference methods. Experimental results demonstrate that MMLC-owf outperforms state-of-the-art methods and MMLC-df enhances the quality of existing truth inference methods.
1 Introduction
@@
Truth inference in crowdsourcing aims to derive accurate results from noisy data provided by online workers. Existing truth inference methods can be broadly classified into two categories: weakly supervised and supervised approaches. In the weakly supervised approach, unknown ground truth is treated as hidden variables. This involves utilizing statistics from workers’ noisy answers to calculate results directly. Alternatively, it entails creating worker behavior models and employing unsupervised learning methods such as the EM algorithm to estimate unknown parameters and infer the ground truth. The weakly supervised approach can further be classified into statistical learning and deep learning methods based on whether it consider the features of the items. Statistical learning methods, such as Majority voting (MV) Imamura et al., (2018), Dawid&Skene model (DS) Dawid and Skene, (1979), and homologous Dawid-Skene model (HDS) Karger et al., (2011); Li and Yu, (2014), do not incorporate item features. In contrast, deep learning methods like Training Deep Neural Nets Gaunt et al., (2016) take item features into account. In the supervised approach, a classifier model is first constructed with item features as the input and ground truth as the output. Then, a worker behavior model is created based on a confusion matrix that establishes the relationship between the item’s ground truth and the worker labels. On this basis, supervised learning is implemented using the classifier model and the worker behavior model by treating the worker labels as supervisory information. Finally, the output of the classifier model is used as the inferred ground truth. In recent years, various truth inference methods based on supervised learning have been proposed, such as CrowdlayerRodrigues and Pereira, (2018), CoNALChu et al., (2021), and UnionNetWei et al., (2022). However, the worker behavior model based on the confusion matrix faces challenges in effectively capturing variations in feature characteristics across different items. Neglecting these variations in worker behavior under different conditions can result in inaccurate representations of worker behavior, consequently impacting the quality of truth inference. For example, in handwritten digit recognition, workers generally have high accuracy. Suppose there are two items: one closely resembles the digit “1,” but its ground truth is actually “7,” and the other is a normal “7.” The former receives many labels as “1,” while the latter rarely gets labeled as “1.” Under the worker behavior model based on the confusion matrix, it is difficult to accurately model the labeling behavior of such high-difficulty items. Therefore, there is a high probability that the model will interpret the former with “1” as the ground truth, leading to incorrect judgments. The quality of truth inference is influenced by uncertainty from hidden variables, the method’s data adaptability, and the accuracy in characterizing worker behavior. The purpose of this paper is to develop a supervised model that can achieve high-quality truth inference with a worker behavior model at the item feature level.
@@
In this paper, we propose Multi-task Supervised Learning from Crowds (MLC), a novel paradigm for crowdsourcing learning. Unlike traditional paradigm, MLC does not rely on the ground truth of the items but instead focuses on understanding the unique behavior of individual workers across different items. When multiple workers handle the same item, they share the item’s features, leading to a multi-task learning paradigm. Within this paradigm, we propose a method called Mixture of Experts-based Multi-task Supervised Learning from Crowds (MMLC). MMLC does not utilize a single classifier but instead creates multiple expert modules. The outputs from these expert modules serve as the bases of the worker spectral space. Each worker is represented by his or her projection vector in the spectral space that characterizes their behavior. The worker behavior model provides a more precise depiction of their behavior across different items by accurately modeling the workers’ behavior on item features. However, it is important to note that the model itself cannot determine the ground truth. To address this limitation, we introduce two truth inference strategies based on MMLC. The first strategy involves analyzing the distribution of workers’ projections in the worker spectral space. We identify the projection of the oracle worker by applying clustering methods, and consider its labels as the ground truth. This approach is referred to as Oracle Worker Finding of MMLC (MMLC-owf). The second strategy leverages the sparsity of crowdsourced data to fill the original dataset with MMLC outputs, generating a new crowdsourcing dataset. Existing truth inference methods can then be applied within this framework, which is called Data Filling of MMLC (MMLC-df). The main contributions are as follows:
@@
  • We introduce a novel paradigm of multi-task supervised learning from crowds and propose a new worker behavior model called MMLC. This model is well-suited for crowdsourcing learning and offers a more accurate way to characterize worker behavior in item labeling.
@@
  • We leverage MMLC to identify the oracle worker for labeling items as the ground truth, referred to as MMLC-owf. Experimental results demonstrate that the labels obtained using this method exhibit higher quality compared to state-of-the-art methodes.
@@
  • We introduce a truth inference framework called MMLC-df, which leverages the MMLC model to fill sparse crowdsourced data. This framework then applies truth inference methods to determine the ground truth. Experimental results demonstrate that MMLC-df significantly enhances truth inference methods, leading to higher-quality results.
2 Related Work
Weakly supervised approach: This class of methods focuses on modeling the relationship between workers’ noisy answers and the ground truth. In this approach, the ground truth is treated as a hidden variable, and weakly supervised learning techniques are employed to infer the ground truth. The most direct and widely used method is Majority Voting (MV) Sheng et al., (2017). It involves counting the responses from workers and considering the answer with the majority of votes as the ground truth. However, MV overlooks the variations in worker behavior during the labeling process and treats all workers’ labels equally. To address this limitation, Tao et al. Tao et al., (2020) proposed a method that considers the deterministic information of majority and minority categories separately. They establish a voting model that takes into account the labeling quality of workers in different situations. The Dawid-Skene (DS) Dawid and Skene, (1979) method utilizes a confusion matrix to describe each worker’s behavior when handling an item. The EM algorithm is then used to estimate the worker’s confusion matrix parameters and the ground truth. HDS Raykar et al., (2010) assumes equal probabilities of wrong worker selections while evolving the confusion matrix. GLAD Whitehill et al., (2009) not only considers workers’ abilities but also accounts for the difficulty of item processing. It creates a worker behavior model using a sigmoid function and employs the EM algorithm to determine the ground truth of an item. Some weakly supervised inference methods utilize deep learning techniques to infer the ground truth. These methods typically use aggregation algorithms to construct an initial answer table, which combines the noisy labels provided by workers. Subsequently, a neural network model is trained and used for classification based on the selected label set. While these methods Gaunt et al., (2016); Ghiassi et al., (2022); Zhu et al., (2023) introduce reliability measurements of labels, the accuracy of these measurements directly affects the results of truth inference. Existing weakly supervised truth inference methods have achieved some degree of success. However, the data provided by workers is often sparse. Additionally, they treat the ground truth as a hidden variable. As a result, the accuracy of truth inference remains limited.
@@ 2 Related Work
Supervised approach: This approach primarily involves building a classifier model that establishes a connection between item features and the corresponding ground truth. Additionally, it constructs a worker behavior model based on the confusion matrix, which captures the relationship between the item’s ground truth and the labels provided by workers. These worker labels are used as supervision information, and the two models are jointly trained in a supervised learning manner. Ultimately, the output of the classifier model is used as the ground truth for inference Chu and Wang, (2021); Ibrahim et al., (2023). Various techniques have been introduced within this paradigm. For instance, the Expectation-Maximization (EM) algorithm has been employed to integrate label aggregation and classifier training for truth inference. Crowdlayer Rodrigues and Pereira, (2018) introduces the concept of a crowd layer, which replaces the traditional confusion matrix used in the DS model. This approach seamlessly integrates label reasoning and classifier training in an end-to-end manner, resulting in more precise and reliable outcomes. Tan and Chen applied the worker’s confusion matrix Tanno et al., (2019) and the labeled transfer matrix Cao et al., (2023) along with classifier predictions to optimize model parameters and improve the accuracy of item processing. Gao et al., (2022) introduces worker weight vectors, while Cao et al., (2023) focuses on worker label reliability modeling. These approaches aim to estimate the characteristics of workers’ abilities. UnionNet Wei et al., (2022) considers all workers collectively and combines their annotations into a parameter transfer matrix. It then aggregates classifier predictions to facilitate model training. The CoNAL method Chu et al., (2021) effectively handles noise from different sources by categorizing labeling noise into common noise and individual noise. Although supervised learning can be used to infer worker parameters, it struggles to accurately characterize worker behavior. Consequently, when the discrimination among candidate answers is poor, it is difficult to achieve good results using this approach.
Diverging from these methods, our method does not treat the ground truth of the item as a latent variable. Instead, we focus on modeling the relationship between the behavior of workers and the difficulty of item processing. This allows us to establish a supervised learning framework to obtain a worker behavior model. Subsequently, we use this model to identify oracle worker who can accurately label items, thereby determining the ground truth. Alternatively, we use this model to fill in sparse crowdsourced data and then employ existing truth inference methods to obtain the ground truth from the completed data.
3 Problem Formulation
@@
This section provides a formal problem description. Our main goal is to create a worker behavior model and achieve joint learning of worker abilities by utilizing multi-task learning to infer the ground truth. Let $\mathcal{W}=\{w_{j}\}$ denote the worker set, where $w_{j}$ represents an individual worker, and $\mathcal{X}=\{x_{i}\}$ denote the set of items, where $x_{i}$ represents a single item to be labeled. The labels for each item belong to the category set $\mathcal{K}=\{k\}$. We use $y_{ij}$ to denote the category label assigned by worker $w_{j}$ to item $x_{i}$. We have an indicator function $I_{ij}$, where $I_{ij}=1$ if $y_{ij}$ exists and $I_{ij}=0$ otherwise. Consequently, we obtain the crowdsourced triples dataset $\mathcal{D}=\{<x_{i},w_{j},y_{ij}>|I_{ij}=1\}$. With regards to truth inference in crowdsourcing, we provide the following definition:
Definition 1 (Problem of Truth Inference (TI Problem))
@@
By modeling and learning from the crowdsourced label dataset $\mathcal{D}$, the problem of truth inference aims to find a function $g^{*}:\mathcal{X}\rightarrow\mathcal{K}$ such that
@@
Here, $\mathcal{H}$ denotes the hypothesis space of functions, $\|\cdot\|_{\mathcal{H}}$ denotes the norm of hypothesis space, $\lambda$ is the regularization coefficient, $\mathcal{L}$ is the loss function, and $\hat{z}_{i}=t_{i}(\mathcal{D})$ is the estimation of label $z_{i}$ for item $x_{i}$ from learning on the dataset. Since the crowdsourced truth inference problem is an unsupervised learning problem without supervised information, the estimation of ground truth is utilized instead of the goal of learning.
Definition 2 (Problem of Multi-task supervise Learning from Crowds (MLC Problem))
Let $\mathcal{S}_{j}=\{(<x_{i},w_{j}>,y_{ij})\}_{x_{i}\in\mathcal{X}_{j}}$ denote the crowdsourced training dataset for worker $w_{j}$, where $\mathcal{X}_{j}=\{x_{i}|I_{ij}=1\}$. The labels provided by worker $j$ can be regarded as the $j$-th task for the corresponding item. Consequently, we obtain the dataset as $\mathcal{S}=\bigcup_{j}\mathcal{S}_{j}$. The problem of multi-task supervised learning from crowds is to find a worker behavior function $f^{*}:\mathcal{X}\times\mathcal{W}\rightarrow\mathcal{K}$ such that
$f^{*}=\arg\min_{f\in\mathcal{H}}\sum_{j=1}^{|\mathcal{W}|}\frac{1}{|\mathcal{X}_{j}|}\sum_{x_{i}\in\mathcal{X}_{j}}\mathcal{L}\left(y_{ij},f(<x_{i},w_{j}>)\right)+\lambda\|f\|_{\mathcal{H}}.$ (2)
We can observe that the solution to MLC problem does not directly address TI problem. Therefore, we provide two approaches to tackle this issue. The first approach is to identify an oracle worker $w_{oracle}$ based on the distribution of workers. We then consider the labels provided by this oracle worker as the ground truth, that is,
$g^{*}(x_{i})=f^{*}(<x_{i},w_{oracle}>).$ (3)
@@
$\mathcal{D}^{\prime}=\mathcal{D}\cup\left\{<x_{i},w_{j},\hat{y}_{ij}>\Big{|}\hat{y}_{ij}=f^{*}(<x_{i},w_{j}>),I_{ij}=0\right\}.$ (4)
4 Proposed Methodology
@@
To address the MLC problem, we propose a Mixture of Experts based Multi-task Supervised Learning from Crowds (MMLC) model. This model utilizes mixture of experts to characterize the varying attention of workers towards different item features, , aiming to capture the feature-level behavior differences of workers when dealing with various items. The framework of the model is shown in Fig. 1. It consists of three main modules: expert module, gate module, and output module.
@@
$f_{\bm{\Theta}}(<x_{i},w_{j}>)=\bm{o}\left(proj_{\bm{U}(x_{i})}(w_{j})\right),$ (6)
@@
$\mathcal{L}_{\bm{\Theta}}=-\frac{1}{|\mathcal{D}|}\sum_{w_{j}\in\mathcal{W}}\sum_{x_{i}\in\mathcal{X}_{j}}\sum_{k\in\mathcal{K}}y_{ij}^{k}\log\left(f_{\bm{\Theta}}(<x_{i},w_{j}>)\right)+\lambda\|\bm{\Theta}\|_{F},$ (7)
The first term denotes the multi-class cross-entropy loss, while the second term represents the regularization of the model’s parameter set $\bm{\Theta}$ to prevent overfitting. In the equation, $\lambda$ is the regularization coefficient, and $\|\cdot\|_{F}$ denotes the Frobenius norm. By minimizing the loss function, we can obtain the final model $\mathcal{M}^{*}:f_{\bm{\Theta}^{*}}(\cdot)$. This model uses the function $f_{\bm{\Theta}^{*}}$ to predict the labels of worker $w_{j}$ for item $x_{i}$, where $\bm{\Theta}^{*}$ represents the optimized parameters.
Truth Inference by Oracle Worker Finding (MMLC-owf): The MMLC model does not directly generate the ground truth for inference. To address this issue, this section proposes a method for inferring the ground truth by identifying the oracle worker’s projection vector in the worker spectral space. Specifically, each worker is theoretically associated with a projection in the worker spectral space, representing their unique characteristics. Workers are distributed in the spectral space. We assume the existence of an omniscient oracle worker who possesses a projection vector in the spectral space and is capable of providing the ground truth in the MMLC model. Therefore, by identifying the projection vector of the oracle worker in the worker spectral space, we can consider its output as the inferred truth. If we treat any worker as a random expression of the oracle worker’s error, then the center of the worker’s distribution projected onto the spectral space can be regarded as the projection vector of the oracle worker, that is,
@@
$f_{\bm{\Theta}^{*}}(<x_{i},w_{oracle}>)=\bm{o}\left(\bm{U}(x_{i})\bm{v}_{oracle}\right).$ (9)
@@
Truth Inference Framework by Data Filling (MMLC-df): In addition to the MMLC-owf method, we propose a truth inference framework using data filling under the MMLC model called MMLC-df, which utilizes the sparsity of crowdsourced data. A new crowdsourced dataset $\mathcal{D}^{\prime}$ is constructed through data filling as follows:
$\mathcal{D}^{\prime}=\mathcal{D}\cup\left\{<x_{i},w_{j},\hat{y}_{ij}>\Big{|}\hat{y}_{ij}=f_{\bm{\Theta}^{*}}(<x_{i},w_{j}>),I_{ij}=0\right\}.$ (10)
@@
Subsequently, any truth inference method applied to this new crowdsourced dataset can infer higher-quality ground truth compared to that obtained from the original data.
5 Experiments
@@
In this section, we verify the effectiveness of our method through experiments. Our MMLC-owf method is a truth inference method, and we compare it with the following baselines: MVSheng et al., (2008) directly uses majority voting to determine the ground truth; DSDawid and Skene, (1979) employs a confusion matrix to characterize the labeling behavior of workers and uses the EM algorithm to infer the ground truth; HDSKarger et al., (2011) simplifies the DS method by assuming that each worker has the same probability of being correct under different truth values and equal probabilities for incorrect options; FDSSinha et al., (2018) is a simple and efficient algorithm based on DS, designed to achieve faster convergence while maintaining the accuracy of truth inference; Max-MIGCao et al., (2019) utilizes the EM algorithm to integrate label aggregation and classifier training; CoNALChu et al., (2021) distinguishes between common noise and individual noise by predicting a joint worker confusion matrix using classifiers; CrowdARCao et al., (2023) estimates worker capability features through classifier prediction and models the reliability of joint worker labels.
@@
Our truth inference framework, MMLC-df, incorporates a core component that performs data filling. We compare it with the following baselines: G_MV Sheng, (2011) utilizes truth inference results from the MV algorithm to evaluate worker ability and assign new labels accordingly; G_IRT Baker et al., (2017) utilizes joint maximum likelihood estimation to estimate parameters of the IRT model, such as worker abilities and item difficulties, and generates new labels based on these parameters; TDG4Crowd Fang et al., (2023) learns the feature distributions of workers and items separately using worker models and item models. An inference component is used to learn the label distribution and generate new labels.
We conduct experiments using three crowdsourced datasets with item features:
  • LableMe Rodrigues and Pereira, (2018); Russell et al., (2008): This dataset consists of 1000 images categorized into 8 classes, with a total of 2547 labels provided by 59 workers. Each image is represented by 8192-dimensional features extracted using a pre-trained VGG-16 model.

### v2

@@ Abstract
Existing learning-from-crowds methods aim to design proper aggregation strategies to infer the unknown true labels from noisy labels provided by crowdsourcing. They treat the ground truth as hidden variables and use statistical or deep learning based worker behavior models to infer the ground truth. However, worker behavior models that rely on ground truth hidden variables overlook workers’ behavior at the item feature level, leading to imprecise characterizations and negatively impacting the quality of learning-from-crowds. This paper proposes a new paradigm of multi-task supervised learning-from-crowds, which eliminates the need for modeling of items’s ground truth in worker behavior models. Within this paradigm, we propose a worker behavior model at the item feature level called Mixture of Experts based Multi-task Supervised Learning-from-Crowds (MMLC), then, two aggregation strategies are proposed within MMLC. The first strategy, named MMLC-owf, utilizes clustering methods in the worker spectral space to identify the projection vector of the oracle worker. Subsequently, the labels generated based on this vector are regarded as the items’s ground truth The second strategy, called MMLC-df, employs the MMLC model to fill the crowdsourced data, which can enhance the effectiveness of existing aggregation strategies . Experimental results demonstrate that MMLC-owf outperforms state-of-the-art methods and MMLC-df enhances the quality of existing learning-from-crowds methods.
Introduction
@@
Existing learning-from-crowdsmethods can be broadly classified into two categories: weakly supervised and supervised approaches. In the weakly supervised approach, unknown ground truth are treated as hidden variables. This involves utilizing statistics from workers’ noisy answers to calculate results directly. Alternatively, it entails creating worker behavior models and employing unsupervised learning methods such as the EM algorithm to estimate unknown parameters and infer the ground truth. The weakly supervised approach can further be classified into statistical learning and deep learning methods based on whether considering the features of items. Statistical learning methods, such as MV (Imamura, Sato, and Sugiyama 2018), DS (Dawid and Skene 1979), and HDS (Karger, Oh, and Shah 2011; Li and Yu 2014), do not incorporate item features. In contrast, deep learning methods like Training Deep Neural Nets (Gaunt, Borsa, and Bachrach 2016) take item features into account. In the supervised approach, a classifier model is first constructed with item features as the input and ground truth as the output. Then, a worker behavior model is created based on a confusion matrix that establishes the relationship between one item’s ground truth and the worker labels. On this basis, supervised learning is implemented using the classifier model and the worker behavior model by treating the worker labels as supervisory information. Finally, the output of the classifier model is used as the inferred ground truth. In recent years, various learning-from-crowds methods based on supervised learning have been proposed, such as Crowdlayer(Rodrigues and Pereira 2018), CoNAL(Chu, Ma, and Wang 2021), and UnionNet(Wei et al. 2022). However, the worker behavior model based on the confusion matrix faces challenges in effectively capturing variations in feature characteristics across different items. Neglecting these variations in worker behavior under different conditions can result in inaccurate representations of worker behavior, consequently impacting the quality of truth inference. For example, in handwritten digit recognition, workers generally have high accuracy. Suppose there are two items: one closely resembles the digit “1,” but its ground truth is actually “7,” and the other is a normal “7.” The former receives many labels as “1,” while the latter rarely gets labeled as “1.” Under the worker behavior model based on the confusion matrix, it is difficult to model the labeling behavior of such difficult items accurately. Therefore, there is a high probability that the model will interpret the former with “1” as the ground truth, leading to incorrect judgments. The quality of aggregation strategies is influenced by uncertainty from hidden variables, the method’s data adaptability, and the accuracy in characterizing worker behavior. The purpose of this paper is to develop a supervised model that can achieve high-quality truth inference with a worker behavior model at the item feature level.
@@
In this paper, we propose Multi-task Supervised Learning-from-crowds (MLC), a novel paradigm for crowdsourcing learning. Unlike the traditional paradigm, MLC does not rely on the ground truth of the items but instead focuses on understanding the unique behavior of individual workers across different items. When multiple workers handle the same item, they share the item’s features, leading to a multi-task learning paradigm. Within this paradigm, we propose a method called Mixture of Experts-based Multi-task Supervised learning-from-crowds (MMLC). MMLC does not utilize a single classifier but instead creates multiple expert modules. The outputs from these expert modules serve as the bases of the worker spectral space. Each worker is represented by his or her projection vector in the spectral space that characterizes their behavior. The worker behavior model provides a more precise depiction of their behavior across different items by accurately modeling the workers’ behavior on item features. However, it is important to note that the model itself cannot determine the ground truth. To address this limitation, we introduce two aggregation strategies based on MMLC. The first strategy involves analyzing the distribution of workers’ projections in the worker spectral space. We identify the projection of the oracle worker by applying clustering methods, and consider its labels as the ground truth. This approach is referred to as Oracle Worker Finding of MMLC (MMLC-owf). The second strategy leverages the sparsity of crowdsourced data to fill the original dataset with MMLC outputs, generating a new crowdsourcing dataset. Existing learning-from-crowds methods can then be applied within this framework, which is called Data Filling of MMLC (MMLC-df). The main contributions are as follows:
@@
  • We introduce a novel paradigm of multi-task supervised learning-from-crowds and propose a novel worker behavior model called MMLC based on feature-level behavior modeling.
@@
  • We leverage MMLC to identify the oracle worker for labeling items as the ground truth, referred to as MMLC-owf. Experimental results demonstrate that the labels obtained using this method exhibit higher quality compared to state-of-the-art methods.
@@
  • We introduce an aggregation framework called MMLC-df, which leverages the MMLC model to fill sparse crowdsourced data. This framework then applies aggregation methods to determine the ground truth. Experimental results demonstrate that MMLC-df significantly enhances learning-from-crowds methods, leading to higher quality results.
Related Work
Weakly supervised approaches: These approaches model the relationship between workers’ noisy responses and the true labels, treating the ground truth as a latent variable and employing weak supervision techniques to deduce it. MV (Sheng et al. 2017) is a commonly used method that assumes the most frequent response as the ground truth, but it fails to account for worker variability. To address this limitation, (Tao, Jiang, and Li 2020) proposed a model that separately considers the majority and minority responses, factoring in the labeling quality of workers. The DS (Dawid and Skene 1979) uses a confusion matrix to characterize worker behavior and estimates parameters with the EM algorithm to infer the ground truth. HDS (Raykar et al. 2010) posits equal chances of erroneous worker choices, refining the confusion matrix with this assumption. GLAD (Whitehill et al. 2009), conversely, factors in both the proficiency of workers and the inherent challenge of tasks, utilizing a sigmoid function to depict worker behavior and the EM algorithm to ascertain the ground truth. Various weakly supervised learning-from-crowds methods integrate deep learning to deduce the ground truth, initiating with strategies that aggregate the noisy labels into an initial answer table. Post this, a neural network is trained for classification, leveraging the curated label set. While these methods (Gaunt, Borsa, and Bachrach 2016; Ghiassi et al. 2022; Zhu, Xue, and Jiang 2023) propose label reliability metrics that significantly influence the outcomes of crowd-based learning. Additionally, (Xu et al. 2024) presents a two-stage method that uses a multi-centroid grouping penalty to incorporate subgroup structures for tasks and workers in inferring the ground truth. While weakly supervised methods have seen successes, they are often hampered by sparse data and the treatment of ground truth as a hidden variable, which limits their accuracy.
@@
Supervised approaches: These methods focus on creating a classifier to link item features with the ground truth and a worker behavior model using a confusion matrix to reflect the relationship between true labels and worker responses. These responses act as supervision, allowing for joint training of both models in a supervised manner, with the classifier’s output serving as inferred ground truth(Chu and Wang 2021; Ibrahim, Nguyen, and Fu 2023). Techniques like the Expectation-Maximization (EM) algorithm are used for label aggregation and classifier training. Crowdlayer(Rodrigues and Pereira 2018) replaces the traditional confusion matrix with a crowd layer, integrating label reasoning and classifier training for more accurate results. Tan and Chen(Tanno et al. 2019) enhance model accuracy by applying confusion and labeled transfer matrices alongside classifier predictions. Other approaches(Gao et al. 2022; Cao et al. 2023) introduce worker weight vectors and focus on modeling label reliability to estimate worker abilities. UnionNet(Wei et al. 2022) aggregates worker annotations into a parameter transfer matrix to facilitate training. The CoNAL method(Chu, Ma, and Wang 2021) categorizes noise into common and individual types, effectively managing diverse labeling noise. Despite its strengths, supervised learning faces challenges in precisely characterizing worker behavior, making it difficult to achieve optimal results when candidate answers are poorly differentiated.
Departing from the traditional latent variable approach to ground truth, our method concentrates on the interplay between worker behavior and item features. This shift enables us to develop a supervised learning framework from the crowd, yielding a nuanced worker behavior model. With this model, we pinpoint the oracle workers capable of precise labeling, thus ascertaining the ground truth.
Problem Formulation
@@
Our main goal is to obtain a worker behavior model and achieve joint learning of worker abilities by utilizing multi-task learning from crowds to aggregate the ground truth. Let $\mathcal{W}=\{w_{j}\}$ denote the worker set, where $w_{j}$ represents an individual worker, and $\mathcal{X}=\{x_{i}\}$ denote the set of items, where $x_{i}$ represents a single item to be labeled. The labels for each item belong to the category set $\mathcal{K}=\{k\}$. We use $y_{ij}$ to denote the category label assigned by worker $w_{j}$ to item $x_{i}$. We have an indicator function $I_{ij}$, where $I_{ij}=1$ if $y_{ij}$ exists and $I_{ij}=0$ otherwise. Consequently, we obtain the crowdsourced triples dataset $\mathcal{D}=\{\langle x_{i},w_{j},y_{ij}\rangle|I_{ij}=1\}$. With regards to learning from crowd in crowdsourcing, we provide the following definition:
Definition 1
@@
(Problem of Learning-from-Crowds (LC Problem) By modeling and learning from the crowdsourced label dataset $\mathcal{D}$, the problem of learning-from-crowds aims to find a function $g^{*}:\mathcal{X}\rightarrow\mathcal{K}$ such that
@@
Here, $\mathcal{H}$ denotes the hypothesis space of functions, $\|\cdot\|_{\mathcal{H}}$ denotes the norm of hypothesis space, $\lambda$ is the regularization coefficient, $\mathcal{L}$ is the loss function, and $\hat{z}_{i}=t_{i}(\mathcal{D})$ is the estimation of label $z_{i}$ for item $x_{i}$ from learning on the dataset. Since the crowdsourced LC problem is an unsupervised learning-from-crowds problem without supervised information, the estimation of ground truth is utilized instead of the goal of learning.
Definition 2
(Problem of Multi-task supervise Learning-from-Crowds (MLC Problem)) Let $\mathcal{S}_{j}=\{(x_{i},y_{ij})\}_{x_{i}\in\mathcal{X}_{j}}$ denote the crowdsourced training dataset for worker $w_{j}$, where $\mathcal{X}_{j}=\{x_{i}|I_{ij}=1\}$. The labels provided by worker $j$ can be regarded as the $j$-th task for the corresponding item. Consequently, we obtain the dataset as $\mathcal{S}=\bigcup_{j}\mathcal{S}_{j}$. The problem of multi-task supervised learning-from-crowds is to find a worker behavior function $f^{*}\in\mathcal{H}$ such that
$f^{*}=\arg\min_{f\in\mathcal{H}}\sum_{w_{j}\in\mathcal{W}}\frac{1}{|\mathcal{X}_{j}|}\sum_{x_{i}\in\mathcal{X}_{j}}\mathcal{L}\left(y_{ij},f_{w_{j}}(x_{i})\right)+\lambda\|f\|_{\mathcal{H}},$ (2)
where $\mathcal{H}$ is a vector-valued reproducing kernel Hilbert space with functions $f:\mathcal{X}\rightarrow\mathcal{K}^{|\mathcal{W}|}$ having components $f_{j}:\mathcal{X}\rightarrow\mathcal{K}$
We can observe that the solution to the MLC problem does not directly address the LC problem. Therefore, we provide two approaches to tackle this issue. The first approach is to identify an oracle worker $w_{oracle}$ based on the distribution of workers. We then consider the labels provided by this oracle worker as the ground truth, that is,
$g^{*}(x_{i})=f_{w_{oracle}}^{*}(x_{i}).$ (3)
@@
$\mathcal{D}^{\prime}=\mathcal{D}\cup\left\{\langle x_{i},w_{j},\hat{y}_{ij}\rangle\Big{|}\hat{y}_{ij}=f_{w_{j}}^{*}(x_{i}),I_{ij}=0\right\}.$ (4)
Proposed Methodology
@@
To address the MLC problem, we propose a Mixture of Experts based Multi-task Supervised Learning-from-Crowds (MMLC) model. This model utilizes mixture of experts to characterize the varying attention of workers towards different item features, aiming to capture the feature-level behavior differences of workers when dealing with various items. The framework of the model is shown in Fig. 1. It consists of three main modules: expert module, gate module, and output module.
@@
Here, the attention layer helps reduce the model’s dependence on unimportant or redundant features, improving the model’s efficiency and accuracy by focusing on the most useful information.
@@
$f_{w_{j}}(x_{i}|\bm{\Theta})=\bm{o}\left(proj_{\bm{U}(x_{i})}(w_{j})\right),$ (6)
@@
$\mathcal{L}_{\bm{\Theta}}=-\frac{1}{|\mathcal{D}|}\sum_{w_{j}\in\mathcal{W}}\sum_{x_{i}\in\mathcal{X}_{j}}\sum_{k\in\mathcal{K}}y_{ij}^{k}\log\left(f_{w_{j}}(x_{i}|\bm{\Theta})\right)+\lambda\|\bm{\Theta}\|_{F},$ (7)
The first term denotes the multi-class cross-entropy loss, while the second term represents the regularization of the model’s parameter set $\bm{\Theta}$ to prevent overfitting. In the equation, $\lambda$ is the regularization coefficient, and $\|\cdot\|_{F}$ denotes the Frobenius norm. By minimizing the loss function, we can obtain the final model $\mathcal{M}^{*}:f(\cdot|\bm{\Theta}^{*})$. This model uses the function $f_{w_{j}}(x_{i}|\bm{\Theta}^{*})$ to predict the labels of worker $w_{j}$ for item $x_{i}$, where $\bm{\Theta}^{*}$ represents the optimized parameters.
MMLC with Oracle Worker Finding (MMLC-owf): The MMLC model does not directly generate the ground truth for inference. To address this issue, this section proposes a method for inferring the ground truth by identifying the oracle worker’s projection vector in the worker spectral space. Specifically, each worker is theoretically associated with a projection in the worker spectral space, representing their unique characteristics. Workers are distributed in the spectral space. We assume the existence of an omniscient oracle worker who possesses a projection vector in the spectral space and is capable of providing the ground truth in the MMLC model. Therefore, by identifying the projection vector of the oracle worker in the worker spectral space, we can consider its output as the inferred truth. If we treat any worker as a random expression of the oracle worker’s error, then the center of the worker’s distribution projected onto the spectral space can be regarded as the projection vector of the oracle worker, that is,
@@
$f_{w_{oracle}}(x_{i}|\bm{\Theta}^{*})=\bm{o}\left(\bm{U}(x_{i})\bm{v}_{oracle}\right).$ (9)
@@
MMLC with Data Filling (MMLC-df): In addition to the MMLC-owf method, we propose a method using data filling under the MMLC model called MMLC-df, which utilizes the sparsity of crowdsourced data. A new crowdsourced dataset $\mathcal{D}^{\prime}$ is constructed through data filling as follows:
$\mathcal{D}^{\prime}=\mathcal{D}\cup\left\{\langle x_{i},w_{j},\hat{y}_{ij}\rangle\Big{|}\hat{y}_{ij}=f_{w_{j}}(x_{i}|\bm{\Theta}^{*}),I_{ij}=0\right\}.$ (10)
@@
Subsequently, any learning-from-crowds method applied to this new crowdsourced dataset can infer higher-quality ground truth compared to that obtained from the original data.
Experiments
@@
We verify the effectiveness of our method through experiments [^2] Our code is available at https://github.com/Crowds24/MMLC.. We compare our learning-from-crowds method MMLC-owf with the following baselines: MV(Sheng, Provost, and Ipeirotis 2008) directly uses majority voting to determine the ground truth; DS(Dawid and Skene 1979) employs a confusion matrix to characterize the labeling behavior of workers and uses the EM algorithm to infer the ground truth; HDS(Karger, Oh, and Shah 2011) simplifies the DS method by assuming that each worker has the same probability of being correct under different truth values and equal probabilities for incorrect options; FDS(Sinha, Rao, and Balasubramanian 2018) is a simple and efficient algorithm based on DS, designed to achieve faster convergence while maintaining the accuracy of truth inference; Max-MIG(Cao et al. 2019) utilizes the EM algorithm to integrate label aggregation and classifier training; CoNAL(Chu, Ma, and Wang 2021) distinguishes between common noise and individual noise by predicting a joint worker confusion matrix using classifiers; CrowdAR(Cao et al. 2023) estimates worker capability features through classifier prediction and models the reliability of joint worker labels.
@@
We compare MMLC-df with the following baselines: G_MV (Sheng 2011) utilizes truth inference results from the MV algorithm to evaluate worker ability and assign new labels accordingly; G_IRT (Baker, Kim et al. 2017) utilizes joint maximum likelihood estimation to estimate parameters of the IRT model, such as worker abilities and item difficulties, and generates new labels based on these parameters; TDG4Crowd (Fang et al. 2023) learns the feature distributions of workers and items separately using worker models and item models. An inference component is used to learn the label distribution and generate new labels.